Questinn Yala nc has conducted the olbiriq arial costs s related b,たs pruduct liries, usiriq a radburial cusuriq svs ern rolurre as d and an activit 七ased costino system ath he traditional and the acti l based costing sys ems ric ude direct materials and direct labor Tutal Custs Sales Traditiona ABC Products Praduct S540* Product 137t Fraduct 2495 3195,200 梟56,000 46,600 55,000 3T,000 10,000 37,400 158,000 84,000 For each product line, comcute operaeng inoome using=ns traditional odstrg srster produに540 Product 137Yt Producz 2495 rcr each produt Ine, comcute ope tng income using the actv-based costng systam Product 540% Product 13/Y ^ Product 2499 Using the rolla ng torn la concube the perce ca e dr erence r ceraur g con e r each of tre produc eso ala pera ng [ncome ADC Operat ne- e trad onal cost ÷ Operating Incore traditiona cost), dem ras ers to 2 dec mal p ces e.g. 1z.25. Enter negative amounts using either a negative sign preceding the number eg, -5 arentheses e.g. (45).) Producz 540X Produc: 137Y Producz 2495 91
Answer #1
a
Product 540X 139000 =195000-56000
Product 137Y 103000 =158000-55000
Product 249S 74000 =84000-10000
b
Product 540X 148400 =195000-46600
Product 137Y 121000 =158000-37000
Product 249S 46600 =84000-37400
c
Product 540X 6.76% =(148400-139000)/139000
Product 137Y 17.48% =(121000-103000)/103000
Product 249S -37.03% =(46600-74000)/74000